OpenStudy (megatron):

Which is a factor of 4x^2y+24x^2-24y-144? A. 3 B. y+4 C. x^2-4 D. y+6

OpenStudy (anonymous):

= 4(x^2y + 6x^2 - 6y - 36) = 4(y(x^2 - 6) + 6(x^2 - 6)) = 4(x^2-6)"(y + 6)" D